If a business gains 8% in one year and loses 3% the following year, what is the percentage gain over two years?
Answer by rfer(16322)

1000*1.08=1080/1.03=1048.54
48.54/1000=0.04854=4.854%
I like to use fictitous figures for these kind of problems.
It makes it easier to see what is happening.